# FastSet: Parallel Claim Settlement

### v1

@@ A.3 Templates and Parameters
A.4 Assumptions and Non-Assumptions about Claim Blocks
@@
For example, in the Dynamic Multi-Sig Account example in Section 4.3.4, an instance claim that is currently valid may become invalid, if the contract lowers its quorum or removes some verifiers. Another example is the voting example in Section 4.9. As explained there, if the voting authority tries to count the number of registered/voted voters and then take according actions that affect the voters’ claims/blocks—such as closing the voting as soon as it collects 1000 votes—it breaks the weakly independence claim assumption. Indeed, suppose the authority currently collects 999 votes. It is still possible and valid to process Alice’s vote, or Bob’s vote, but not both of them.

### v3

@@ A.3 Templates and Parameters
A.4 Claim Block Assumptions and Non-Assumptions
@@
For example, in the Dynamic Multi-Sig Account example in Section 4.3.4, an instance claim that is currently valid may become invalid if the contract lowers its quorum or removes some verifiers. Another example is the voting example in Section 4.9. As explained there, if the voting authority tries to count the number of registered/voted voters and then take according actions that affect the voters’ claims/blocks—such as closing the voting as soon as it collects 1000 votes—it breaks the weakly independence claim assumption. Indeed, suppose the authority currently collects 999 votes. It is still possible and valid to process Alice’s vote, or Bob’s vote, but not both of them.